# AutoGCN - Towards Generic Human Activity Recognition with Neural Architecture Search

### 02 Feb 2024 v1

@@ Abstract
This paper introduces AutoGCN, a generic Neural Architecture Search (NAS) algorithm for Human Activity Recognition (HAR) using Graph Convolution Networks (GCNs). HAR has gained attention due to advances in deep learning, increased data availability, and enhanced computational capabilities. At the same time, GCNs have shown promising results in modeling relationships between body key points in a skeletal graph. While domain experts often craft dataset-specific GCN-based methods, their applicability beyond this specific context is severely limited. AutoGCN seeks to address this limitation by simultaneously searching for the ideal hyperparameters and architecture combination within a versatile search space using a reinforcement controller while balancing optimal exploration and exploitation behavior with a knowledge reservoir during the search process. We conduct extensive experiments on two large-scale datasets focused on skeleton-based action recognition to assess the proposed algorithm’s performance. Our experimental results underscore the effectiveness of AutoGCN in constructing optimal GCN architectures for HAR, outperforming conventional NAS and GCN methods, as well as random search. These findings highlight the significance of a diverse search space and an expressive input representation to enhance the network performance and generalizability.
@@ 1 Introduction
Human Activity Recognition (HAR), the process of identifying and categorizing human activities, has witnessed a surge in interest in recent years. This momentum can be attributed to advancements in deep learning techniques, the increased availability of data, and enhanced computational capabilities [1]. HAR finds its applications in various domains such as video surveillance, human-computer interaction, and healthcare, highlighting its significance and broad impact [1, 2, 3, 4]. Within the realm of HAR, the primary models aim to capture representations of the human with a skeleton graph built of specific body key points [1]. In the context of HAR, the traditional feature extraction method relies on manually crafted descriptors to capture fundamental attributes [5, 6]. Another technique involves conceptualizing the problem as a deep learning task within Euclidean space. This approach can then use the advanced capabilities of Convolutional Neural Networks and Recurrent Neural Networks associated with recognition tasks. Yet, approaches involving these architectures have a significant shortcoming in modeling the inherent structures in skeletons since the input data gets serialized [1]. Recently, Graph Neural Networks (GNNs), specifically Graph Convolution Networks (GCNs), have emerged as a cutting-edge solution in HAR [7]. GCNs have achieved remarkable performance by leveraging the relationships between different segments and joints, treating these as nodes and edges in an adjacency matrix, making convolutional operations possible [1, 7].
While GCN-based methods have shown promising results, they are primarily developed by domain experts and often tailored to specific datasets through a trial-and-error process. It can be challenging to apply methods to a broader range of problems or when domain knowledge is unavailable, limiting their potential impact. One approach to overcome this challenge is automatically building task-specific deep learning models using Neural Architecture Search (NAS) [8]. By employing NAS, researchers can discover optimal architectures from a defined search space through a search procedure, enhancing the overall performance and generalizability of the resulting models.
In this work, we propose a generic NAS algorithm, AutoGCN, that dynamically learns both hyper- and architecture parameters to construct a framework that maximally fits the HAR task. This paper makes the following contributions:
  • A generic approach to construct a GCN architecture for HAR with NAS, which can concurrently optimize hyperparameters and the architecture. Additionally, it can be applied to different datasets within HAR and performs better than random search and other NAS procedures for GCNs.
@@
  • A diverse search space construction enables the usage of different input representations further to enhance the network performance and generalizability.
@@ 2.1 Skeleton-based HAR
HAR within the domain of skeletal data attracts increasing interest due to their rectified robustness against background noise, occlusion, or scaling implications when compared to conventional RGB methods [3]. Since the nature of the skeleton data modality is non-euclidean, current approaches commonly use GCNs, as these can operate on unstructured data. The first work in this field stems from [7], introducing ST-GCN to learn spatial and temporal patterns from the input data with a tailored GCN layer. This foundation makes the handcrafted feature engineering obsolete while achieving superior performances compared to previous methods [7]. Several subsequent studies have adopted this technique and introduced multi-stream architectures to incorporate different kinematic variables from which the model can learn. A two-stream adaptive graph convolutional network named 2s-AGCN was introduced in [9] to consider the skeleton’s second-order information and learn the graph’s topology. In a later study [10], Efficient-GCN was introduced, built upon three input branches and incorporating techniques from separable convolution to make the network more efficient. While all of these approaches achieve good performances, they are all designed by human experts, hindering their generalizability to scenarios outside the tested databases and development by inexperienced practitioners. This leaves the automatic construction of generic models an open problem.
@@ 2.2 Neural Architecture Search
For most tasks, deep neural networks must have a customized architecture to accomplish their objectives effectively. Initially, using deep neural networks automated the manual feature engineering aspect of machine learning. The field has since progressed further with the introduction of NAS and automated machine learning (AutoML), which automates the construction of neural architectures and accompanying tasks like hyperparameter optimization, e.g., batch size, learning rate, or algorithm selection [11]. These approaches have already given rise to architectures in the area of image classification that outperform models designed by human experts, as highlighted in studies such as [8, 12].
@@
A new approach named AutoHAS incorporating hyperparameters and architecture components in the search was introduced in [19]. This comprehensive approach aims to optimize the architectural design and the associated hyperparameters like the chosen optimizer and its learning rate for increased performance. To accomplish this, the shared network weights and a reinforcement learning controller are updated alternately to find the best probability distribution in the combined search space, resulting in the best-performing architecture and hyperparameter configuration. By extending the framework to incorporate hyperparameter optimization, AutoHAS went beyond the traditional scope of NAS and encompassed a broader spectrum of optimization possibilities. The inclusion of hyperparameter search is a significant enhancement as it allows for the tuning of various factors that impact model performance, such as learning rates, regularization parameters, and batch sizes, which have a significant impact on the model performance [12]. In contrast, AutoHAS relies on the straightforward Reinforce algorithm [20], which overlooks previously sampled and trained candidate architectures during the update procedure, resulting in an imbalance between exploration and exploitation behaviors. Additionally, the architectural search is confined to basic building blocks, and the exploration of hyperparameters is restricted to selecting the most suitable learning rate and optimizer.
@@
To date, the research of NAS has primarily been focused on and applied to image classification tasks. In the domain of HAR GCNs, the NAS research focus has been limited to [21, 22], to the best of our knowledge. In the work of [21], NAS was employed to discover the most effective architecture by approximating spatiotemporal cues and leveraging Chebyshev polynomials of varying orders. This approach involved constructing a search space consisting of three dynamic graph substructures: spatial, temporal, and spatiotemporal. Additionally, the authors searched for eight function modules that were subsequently applied to each network layer. Expanding upon these concepts, SNAS-GCN represents an extension of this research, optimizing the search space and implementing a single-path one-shot approach for improved search efficiency [22]. The experimental findings of SNAS-GCN demonstrate a reduction in search time compared to the previous work [21], albeit with a diminished level of accuracy. Despite the significant contributions of [21, 22] to the field of skeleton GCNs through utilizing NAS, their work has certain limitations. One limitation lies in their reliance on approximating spatiotemporal cues and utilizing Chebyshev polynomials, which can lead to information loss and impact the derived architectures’ accuracy and robustness. Furthermore, the search for a modest set of eight function modules applied to the network layers may restrict the expressiveness and flexibility of the discovered architectures. The lack of flexibility in the network architecture is a further notable limitation, as it is fixed to ten layers with uniform channel sizes and composition. This rigidity restricts the expressiveness of the architecture, potentially hindering its ability to effectively capture complex patterns in the data. In addition, the authors only search for the spatial and temporal filters solely on the joint data while sharing them with the other four dataset modalities within NTU RBG+D [23], and hereby overlook the potential variance present in these modalities. Moreover, the final performance of the architecture relies on an ensemble score derived from two distinct data modalities, making it necessary to use more resources to search and train those models. Considering these limitations, a more extensive and diverse search space combined with an ample data representation, the model could leverage a richer set of operations and adapt better to the specific task requirements, potentially leading to improved performance.
@@ 3 AutoGCN
AutoGCN is a novel NAS framework that predicts action classes for a given sequence of skeletons. In this section, we provide AutoGCN’s technical details. First, we introduce some preliminaries and the data representation as the foundational input features. Second, we discuss the defined search space and its implications. Third and finally, we explain the search algorithm in detail. Fig. 1 illustrates the workflow in the AutoGCN framework.
@@ 3.2 Data representation
Prior research on skeleton-based action recognition has demonstrated the paramount importance of data preprocessing [24, 25, 26, 27, 9, 28, 29]. In the context of this study, the input features can be categorized into four distinct groups, namely: 1) position $P$, 2) velocity $V$, 3) bone $B\in\{L,\beta\}$, and 4) acceleration features $A$, corresponding to the fundamental kinematic properties of the human body [30]. The input sequence $\mathcal{X}$, denoted as $\mathcal{X}\in\mathcal{R}^{C_{in}\times T_{in}\times V_{in}\times M_{in}}$ represents a sequence of skeleton data inputs, where $C_{\text{in}}$ denotes the number of channels indicating the $\{x,y,z\}$ orientation, $T_{\text{in}}$ represents the number of frames in the sequence, $V_{\text{in}}$ specifies the number of vertices representing spatial points, and $M_{\text{in}}$ depicts the number of skeletons present in the sequence. This input sequence undergoes a subsequent partitioning into the mentioned categories. The relative position $P$ is calculated via
@@
where $c$ is the chosen central joint. The calculation of velocity $V$ is determined via
@@
The parameter $d$ determines the timeframes over which the velocity is measured, with different values yielding different temporal spans. In this study, the range of $d$ is within $[1,2]$. To calculate the acceleration $A$, the velocity from 2 is differentiated with respect to time. Since the equation involves discrete time steps, finite differences can approximate the derivative. The acceleration $A$ is then calculated as follows
@@
Additionally, an 8th-order Butterworth filter is also introduced to compensate by the derivative [31] from the acceleration feature. This meticulous filtering process is critical in enhancing the signal quality, enabling a more accurate and reliable representation of the underlying data. The bone features $B$ are divided into bone length $L$ and respective angles $\beta$ which has been introduced in [10]:
@@ 3.3 Spatial Graph Convolution
Here, $v_{ti}$ represents the $i$-th joint at the $t$-th frame, and $f_{out}(.)$ and $f_{in}(.)$ correspond to the output and input features of the respective joints. Furthermore, $B(v_{ti})$ denotes the set of neighboring joints of $v_{ti}$, while $B_{v_{ti}}$ defines the convolutional sampling area, specifically encompassing vertexes that are $1$-distance neighbors (i.e. immediate neighboring joints). The normalization factor $Z_{ti}(v_{tj})$ represents the size of the corresponding subset, and its purpose is to equalize the influence of various subsets on the final output. The weighting function $\mathbf{w(.)}$ incorporates the mapping function $l_{ti}(.)$, which serves to establish diverse partitioning strategies, such as uni-labeling, distance, and spatial configuration partitioning [7]. To further refine the formulation, Equation 7 is written as
@@ 3.4 Search Space
where the hyperparameter $h$ is a sum of weighted components $C_{i}^{h}$ of the predefined values $\mathcal{B}_{i}$, satisfying two conditions: the weights sum up to 1, to ensure those are normalized and that the combined influence of all components adds up to the final hyperparameter value, and $C_{i}^{h}$ behaves either as a continuous value in $[0,1]$ or a binary choice {0,1} depending on whether it represents a continuous or categorical hyperparameter.
@@
where $n$ defines the chosen architecture type $C_{i,j}^{\alpha}$ and $k$ the corresponding value out of the predefined set of architectural parameters $\Delta_{i,j}$. Just as with the hyperparameters, the architecture can be categorical or continuous. By considering both the architectural components and the hyperparameters, AutoGCN ensures a comprehensive search process, allowing for the fine-tuning of multiple parts to achieve an optimal outcome. Table 1 shows the entire search space and its variables.
@@ 3.5 Architecture
As previously discussed, the architecture construction necessitates careful consideration of the unique data representation intrinsic to the skeleton sequence. To address this requirement effectively, the architecture is thoughtfully partitioned into separate building blocks, building upon prior research [10, 29, 37]. The fundamental elements comprise 1) the initialization layer, 2) the input stream- and, 3) the main stream search block, followed by 4) the classifier block, as depicted in Fig. 2.
@@
Each block integrates various parameters derived from the comprehensive search space in Table 1. This modular approach ensures the architecture can deal with the intricate nuances and complexities inherent in the skeleton sequence data representation.
@@ 3.6 Search Algorithm
AutoGCN seeks optimization on two fronts: the hyperparameters denoted as $h$, and the architecture represented as $\alpha$. This approach aligns with prior research in this field [19, 12]. The overarching objective is to minimize the loss function $\mathcal{L}$ on the validation dataset $\mathcal{D}_{val}$ while ensuring that the optimal parameters $\omega_{\alpha,h}^{*}$ are derived from the minimized loss on both the architecture $\alpha$, hyperparameters $h$ and the training dataset $\mathcal{D}_{train}$. This can be formally articulated as follows:
@@
To address the potential inefficiencies of the Reinforce algorithm’s sampling behavior, AutoGCN incorporates a replay memory $\mathbb{N}$ that stores a collection of $\mathcal{N}$ previously generated student architectures. This memory is then used to enhance the sampling procedure for updating the controller. The addition of a replay memory helps mitigate the limitations of the Reinforce algorithm by utilizing a broader range of student architectures that have been saved over time. This has the effect of expanding the effective sampling size and improving the controller’s decision-making process. Refer to Algorithm 1 for further details.

### 12 Mar 2024 v3

@@ Abstract
This paper introduces AutoGCN, a generic Neural Architecture Search (NAS) algorithm for Human Activity Recognition (HAR) using Graph Convolution Networks (GCNs). HAR has enjoyed increased attention due to advances in deep learning, increased data availability, and enhanced computational capabilities. Concurrently, GCNs have shown promising abilities in modeling relationships between body key points in a skeletal graph. Typically, domain experts develop dataset-specific GCN-based methods, which limits their applicability beyond the specific context. AutoGCN seeks to address this limitation by simultaneously searching for the ideal hyperparameters and architecture combination within a versatile search space using a reinforcement controller while balancing optimal exploration and exploitation behavior with a knowledge reservoir during the search process. We conduct extensive experiments on two large datasets focused on skeleton-based action recognition to assess the proposed algorithm’s performance. Our experimental results demonstrate the effectiveness of AutoGCN in constructing optimal GCN architectures for HAR, outperforming conventional NAS and GCN methods, as well as random search. These findings highlight the significance of a diverse search space and an expressive input representation to achieve good model performance and generalizability.
@@ 1 Introduction
Human Activity Recognition (HAR), the process of identifying and categorizing human activities, has witnessed a surge in interest in recent years. This momentum is attributed to advancements in deep learning techniques, the increased availability of data, and enhanced computational capabilities [1]. HAR has applications in a wide range of domains, including video surveillance, human-computer interaction, and healthcare, highlighting its significance and broad impact [1, 2, 3, 4]. Within the realm of HAR, the models aim to capture representations of humans using a skeleton graph built of body key points [1]. In the context of HAR, the traditional feature extraction method relies on manually crafted descriptors to capture fundamental attributes [5, 6]. Another approach involves formulating the problem as a deep learning task in Euclidean space. This approach uses the capabilities of Convolutional Neural Networks and Recurrent Neural Networks associated with recognition tasks. However, approaches involving these architectures have a significant shortcoming in modeling the inherent structures in skeletons, as the input data is serialized [1]. Recently, Graph Neural Networks (GNNs), specifically Graph Convolution Networks (GCNs), have emerged as a cutting-edge solution in HAR [7]. GCNs have achieved remarkable performance by leveraging the relationships between different segments and joints, treating these as nodes and edges in an adjacency matrix, making convolutional operations possible [1, 7].
While GCN-based methods have shown promising results, they are primarily developed by domain experts and often tailored to specific datasets through a trial-and-error process. This limits their potential impact, as applying them to a broader range of problems can be challenging or when domain knowledge is unavailable. One approach to overcoming those challenges is automatically building task-specific deep learning models using Neural Architecture Search (NAS) [8]. By employing NAS, researchers can discover optimal architectures from a defined search space through a search procedure, enhancing the overall performance and generalizability of the resulting models.
In this work, we propose a generic NAS algorithm, AutoGCN, that dynamically learns both hyper- and architecture parameters to construct a framework that maximally fits the HAR task. We present the following contributions:
  • A generic approach to construct a GCN architecture for HAR with NAS, which can concurrently optimize hyperparameters and the model architecture. Additionally, the approach can be applied to different datasets within HAR and performs better than random search and other NAS procedures for GCNs.
@@
  • A diverse search space construction that enables the usage of different input representations to enhance the network performance and generalizability further.
@@ 2.1 Skeleton-based HAR
HAR within the domain of skeletal data attracts increasing interest due to its rectified robustness against background noise, occlusion, or scaling implications when compared to conventional RGB methods [3]. Since the nature of the skeleton data modality is non-euclidean, current approaches commonly use GCNs, as these can operate on unstructured data. The first work in this field stems from [7], introducing ST-GCN to learn spatial and temporal patterns from the input data with a tailored GCN layer. This foundation makes the handcrafted feature engineering obsolete while achieving superior performances compared to previous methods [7]. Several subsequent studies have adopted this technique and introduced multi-stream architectures to incorporate different kinematic variables from which the model can learn. A two-stream adaptive graph convolutional network named 2s-AGCN was introduced in [9] to consider the skeleton’s second-order information and learn the graph’s topology. In a later study [10], Efficient-GCN was introduced, built upon three input branches and incorporating techniques from separable convolution to make the network more efficient. While these approaches yield good performances, they are all designed by human experts, limiting their generalizability to scenarios beyond the tested databases or usage by non-domain experts. This leaves the problem of the automatic construction of generic models unresolved.
@@ 2.2 Neural Architecture Search
For most tasks, deep neural networks must have a customized architecture to accomplish their objectives effectively. Initially, deep neural networks were used to automate the manual feature engineering aspect of machine learning. The field has since progressed further with the introduction of NAS and automated machine learning (AutoML), which automates the construction of neural architectures and accompanying tasks like hyperparameter optimization, e.g., batch size, learning rate, or algorithm selection [11]. These approaches have already given rise to architectures in the area of image classification that outperform models designed by human experts, as highlighted in studies such as [8, 12].
@@
A new approach named AutoHAS incorporating hyperparameters and architecture components in the search was introduced in [19]. This comprehensive method aims to optimize the architectural design and the associated hyperparameters like the chosen optimizer and its learning rate for increased performance. Therefore, the shared network weights and a reinforcement learning controller are updated alternately to find the best probability distribution in the combined search space, resulting in the best-performing architecture and hyperparameter configuration. By extending the framework to incorporate hyperparameter optimization, AutoHAS went beyond the traditional scope of NAS and encompassed a broader spectrum of optimization possibilities. The inclusion of hyperparameter search is a significant enhancement as it allows for the tuning of various factors that impact model performance, such as learning rates, regularization parameters, and batch sizes, which have a significant impact on the model performance [12]. In contrast, AutoHAS relies on the straightforward Reinforce algorithm [20], which overlooks previously sampled and trained candidate architectures during the update procedure, resulting in an imbalance between exploration and exploitation behaviors. Additionally, the architectural search is confined to basic building blocks, and the exploration of hyperparameters is restricted to selecting the most suitable learning rate and optimizer.
@@
To date, NAS research has primarily focused on and applied to image classification tasks. In the domain of HAR GCNs, the NAS research focus has been limited to [21, 22], to the best of our knowledge. In the work of [21], NAS was employed to discover the most effective architecture by approximating spatiotemporal cues and leveraging Chebyshev polynomials of varying orders. This approach involved constructing a search space consisting of three dynamic graph substructures: spatial, temporal, and spatiotemporal. Additionally, the authors searched for eight function modules that were subsequently applied to each network layer. Expanding upon these concepts, SNAS-GCN represents an extension of this research, optimizing the search space and implementing a single-path one-shot approach for improved search efficiency [22]. The experimental findings of SNAS-GCN demonstrate a reduction in search time compared to the previous work [21], albeit with a diminished level of accuracy. Despite the significant contributions of [21, 22] to the field of skeleton GCNs through utilizing NAS, their work has certain limitations. One limitation lies in their reliance on approximating spatiotemporal cues and utilizing Chebyshev polynomials, which can lead to information loss and impact the derived architectures’ accuracy and robustness. Furthermore, the search for a modest set of eight function modules applied to the network layers may restrict the expressiveness and flexibility of the discovered architectures. The lack of flexibility in the network architecture is a further notable limitation, as it is fixed to ten layers with uniform channel sizes and composition. This rigidity restricts the expressiveness of the architecture, potentially hindering its ability to capture complex patterns in the data effectively. In addition, the authors only search for the spatial and temporal filters solely on the joint data while sharing them with the other four dataset modalities within NTU RBG+D [23], and, as a result of this, overlook the potential variance present in these modalities. Moreover, the final performance of the architecture relies on an ensemble score derived from two distinct data modalities, making it necessary to use more resources to search and train those models. Considering these limitations, a more extensive and diverse search space combined with an ample data representation, the model could leverage a richer set of operations and adapt better to the specific task requirements, potentially leading to improved performance.
@@ 3 AutoGCN
AutoGCN is a novel NAS framework that predicts action classes for a given sequence of skeletons. In this section, we provide AutoGCN’s technical details. First, we introduce some preliminaries and the data representation as the foundational input features. Second, we discuss the defined search space and its implications. Third and finally, we explain the search algorithm in detail. Fig. (1) illustrates the workflow in the AutoGCN framework.
@@ 3.2 Data representation
Prior research on skeleton-based action recognition demonstrated the paramount importance of data preprocessing [24, 25, 26, 27, 9, 28, 29]. In the context of this study, the input features are categorized into four distinct groups, namely: 1) position $P$, 2) velocity $V$, 3) bone $B\in\{L,\beta\}$, and 4) acceleration features $A$, corresponding to the fundamental kinematic properties of the human body [30]. The input sequence $\mathcal{X}$, denoted as $\mathcal{X}\in\mathcal{R}^{C_{in}\times T_{in}\times V_{in}\times M_{in}}$ represents a sequence of skeleton data inputs, where $C_{\text{in}}$ denotes the number of channels indicating the $\{x,y,z\}$ orientation, $T_{\text{in}}$ represents the number of frames in the sequence, $V_{\text{in}}$ specifies the number of vertices representing spatial points, and $M_{\text{in}}$ depicts the number of skeletons present in the sequence. This input sequence undergoes a subsequent partitioning into the mentioned categories. The relative position $P$ is calculated via
@@
where $c$ is the chosen central joint. The velocity $V$ is determined via
@@
The parameter $d$ determines the timeframes over which the velocity is measured, with different values yielding different temporal spans. In this study, the range of $d$ is within $[1,2]$. To calculate the acceleration $A$, the velocity from (2) is differentiated with respect to time. Since the equation involves discrete time steps, finite differences are used to approximate the derivative. The acceleration $A$ is then calculated as follows
@@
An 8th-order Butterworth filter is also introduced to compensate by the derivative [31] from the acceleration feature. This meticulous filtering process is critical in enhancing the signal quality, enabling a more accurate and reliable representation of the underlying data. The bone features $B$ are divided into bone length $L$ and respective angles $\beta$ which have been introduced in [10]:
@@ 3.3 Spatial Graph Convolution
Here, $v_{ti}$ represents the $i$-th joint at the $t$-th frame, and $f_{out}(.)$ and $f_{in}(.)$ correspond to the output and input features of the respective joints. Furthermore, $B(v_{ti})$ denotes the set of neighboring joints of $v_{ti}$, while $B_{v_{ti}}$ defines the convolutional sampling area, specifically encompassing vertexes that are $1$-distance neighbors (i.e., immediate neighboring joints). The normalization factor $Z_{ti}(v_{tj})$ represents the size of the corresponding subset, and its purpose is to equalize the influence of various subsets on the final output. The weighting function $\mathbf{w(.)}$ incorporates the mapping function $l_{ti}(.)$, which serves to establish diverse partitioning strategies, such as uni-labeling, distance, and spatial configuration partitioning [7]. To further refine the formulation, Equation (7) is written as
@@ 3.4 Search Space
where the hyperparameter $h$ is a sum of weighted components $C_{i}^{h}$ of the predefined values $\mathcal{B}_{i}$, satisfying two conditions: the weights sum up to 1, to ensure these are normalized and that the combined influence of all components adds up to the final hyperparameter value, and $C_{i}^{h}$ behaves either as a continuous value in $[0,1]$ or as a binary choice {0,1} depending on whether it represents a continuous or categorical hyperparameter.
@@
where $n$ defines the chosen architecture type $C_{i,j}^{\alpha}$ and $k$ the corresponding value out of the predefined set of architectural parameters $\Delta_{i,j}$. Just as with the hyperparameters, the architecture can be categorical or continuous. By considering both the architectural components and the hyperparameters, AutoGCN ensures a comprehensive search process, allowing for the fine-tuning of multiple parts to achieve an optimal outcome. Table (1) shows the entire search space and its variables.
@@ 3.5 Architecture
As previously discussed, the architecture construction necessitates careful consideration of the unique data representation intrinsic to the skeleton sequence. To address this requirement effectively, the architecture is thoughtfully partitioned into separate building blocks, building upon prior research [10, 29, 37]. The fundamental elements comprise 1) the initialization layer, 2) the input stream- and, 3) the main stream search block, followed by 4) the classifier block, as depicted in Fig. (2).
@@
Each block integrates various parameters derived from the comprehensive search space in Table (1). This modular approach ensures that the architecture can deal with the intricate nuances and complexities inherent in the skeleton sequence data representation.
@@ 3.6 Search Algorithm
AutoGCN seeks optimization on two fronts: the hyperparameters denoted as $h$, and the architecture represented as $\alpha$. This approach aligns with prior research in this field [19, 12]. The overarching objective is to minimize the loss function $\mathcal{L}$ on the validation dataset $\mathcal{D}_{val}$ while ensuring that the optimal parameters $\omega_{\alpha,h}^{*}$ are derived from the minimized loss on both the architecture $\alpha$, hyperparameters $h$ and the training dataset $\mathcal{D}_{train}$. This is articulated as follows:
@@
To address the potential inefficiencies of the Reinforce algorithm’s sampling behavior, AutoGCN incorporates a replay memory $\mathbb{N}$ that stores a collection of $\mathcal{N}$ previously generated student architectures. This memory is used to enhance the sampling procedure for updating the controller. The addition of a replay memory helps mitigate the limitations of the Reinforce algorithm by utilizing a broader range of student architectures that have been saved over time. This has the effect of expanding the effective sampling size and improving the controller’s decision-making process. Refer to Algorithm (1) for further details.